The model airplane has a mass of 0.90 kg and moves at constant speed on a circle that is parallel to the ground.
The path of the airplane and the guideline lie in the same horizontal plane because the weight of the plane is balanced by the lift generated by its wings. Find the tension in the 17 m guideline for a speed of 19 m/s.
Answer #1

T is producing centripetal force

so,

T = Mv^2 / r

T = 0.9 * 19^2 / 17

= 19.11 N <============ Answeer
Answer #2

T = Mv^2 / r

T = 0.9 * 19^2 / 17

= 19.11 N